OpenStudy (anonymous):

find the given zeros of the function f(x)=x^3+3x^2-6x-18 f(x)=x^3+3x^2+6x-18 f(x)=x^3+3x^2-6x+18

ganeshie8 (ganeshie8):

factor first two terms and factor last two terms

OpenStudy (anonymous):

i know, i do but i get like plus or minus negative 6 so do i pull the negative out or?

ganeshie8 (ganeshie8):

f(x)=x^3+3x^2-6x-18 ------- ------ = x^2(x+3) - 6(x+3)

ganeshie8 (ganeshie8):

^^yess

ganeshie8 (ganeshie8):

next factor the gcf again

ganeshie8 (ganeshie8):

f(x)=x^3+3x^2-6x-18 ------- ------ = x^2(x+3) - 6(x+3) = (x+3)(x^2-6)

ganeshie8 (ganeshie8):

you need to factor x^2-6 still

ganeshie8 (ganeshie8):

u may use the identity : a^2-b^2 = (a+b)(a-b)

OpenStudy (anonymous):

so @ganeshie8 which one would give me the the given zeroes of the square root of 6, the negative square root of 6, and -3???

ganeshie8 (ganeshie8):

x^2-6 = x^2 - sqrt(6)^2 = (x+sqrt(6))(x-sqrt(6))

ganeshie8 (ganeshie8):

plug that in, f(x)=x^3+3x^2-6x-18 ------- ------ = x^2(x+3) - 6(x+3) = (x+3)(x^2-6) = (x+3) (x+sqrt(6))(x-sqrt(6))

OpenStudy (anonymous):

thank you!